OpenStudy (anonymous):

Solve each system by elimination -x+5y=-11 x+y=5

OpenStudy (anonymous):

if you add them you get 6y=-6 y=-1 substitute back in to get x-1=5 x=6 so your solution set is (6,-1)

OpenStudy (anonymous):

basically in elimination you want to get to get a negative and positive number of x or y in which when you add, you'll cancel one variable out leaving one left for an easily solve

OpenStudy (anonymous):

Thanks, Guy. I have a similar one

OpenStudy (anonymous):

-4x+y=-4 -2x+4y-16